OpenStudy (anonymous):

write the slope intercept equation for the line that passes thru (-2,3) and perpendicular to 2x-9y=-12. show work

OpenStudy (anonymous):

ok, so the slope intercept equation is y=mx + b

OpenStudy (anonymous):

but first you should isolate y in the equation you already have. So you can isolate the rate of change aka the mx part.

OpenStudy (anonymous):

-9y = -2x - 12 then: y = (2/9)x - 12/9 to simplify: y =(2/9)x - 4/3 aka 1 1/3

OpenStudy (anonymous):

so then, as its perpendicular, you'll know that the slope of the opposite line is the negative recipricol. in this case, -9/2x

OpenStudy (anonymous):

then, input (-2,3) respectfully in the x and y places in the now modified equation. so you get: 3 = (-9/2)(-2) + b so: 3 = 9 + b then: -6 = b

OpenStudy (anonymous):

so you end up with y=-9/2x - 6

OpenStudy (anonymous):

Hope that helped. medal, s'il vous plait? :)